OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Glen decides he only wants caramels and candy corn. Caramels sell for $0.95 per pound, and candy corn sells for $0.79 per pound. Glen’s bag weighs 1.7 pounds and it costs $1.48. If p is the number of pounds of caramels Glen purchases, what expression could be used to represent the number of pounds of candy corn he purchases? (Points : 1) 1.7p 1.7 – p p + 1.7 p – 1.7

OpenStudy (anonymous):

@rebecca99 @Rockiee @Ashleyisakitty @acxbox22 @ganeshie8

OpenStudy (anonymous):

Hello, Do you have any idea how to start this problem or are you completely lost?

OpenStudy (anonymous):

not really i think you take out what you know that they saying which is 1.7 and x

OpenStudy (anonymous):

and hi :)

OpenStudy (anonymous):

Okay, I will try to walk you step by step so you understand the problem. Just tell me if you get lost and where you get lost okay, do not be shy, ask if needed. :-) Caramels = $0.95 per pound You can write the word phrase as a mathematical expression like: \[0.95p~where~p~represents~the~pounds~of~caramels.\] Recall the information given: "If p is the number of pounds of caramels Glen purchases, what expression could be used to represent the number of pounds of candy corn he purchases?" We need the expression that can help us find the pounds of candy corn purchased. We know that the total of the candy purchased weighed \(1.7\) pounds. Since \(p\) represents the pounds of caramels we can subtract \(p\) from \(1.7\) pounds to find the weight of the candy corn.

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Close, however it is not answer choice \(D\). The net weight of the candies is \(1.7\) we need to subtract \(p\) from that rather than \(p\) being subtracted, since \(p\) is not the net aka total weight.

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Correct, congratulations.

OpenStudy (anonymous):

thanks a little more thank you thank you

OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Troy decides he only wants chocolate-covered pretzels and gumdrops. Chocolate-covered pretzels sell for $0.89 per pound, and gumdrops sell for $0.65 per pound. Troy’s bag weighs 1.8 pounds and it cost $1.29. Write an equation in simplest form that represents the problem described here. (Points : 1) 0.24p + 1.17 = 1.8 0.24p + 1.17 = 1.29 –0.24p + 1.602 = 1.29 –0.24p + 1.602 = 1.8

OpenStudy (anonymous):

if you want to help

OpenStudy (anonymous):

OpenStudy has a policy of one question per post normally, close this question and open a new question and I will see if I can help you. Peace.

OpenStudy (anonymous):

okay thanks